Problem 3

At the begening of 2018, a town has 152,000 employed people and 8000 unemployed people. The remaining 40,000

people in the town are not in the labor force.

Based on these​ numbers:

  1. Calculate the​ town's unemployment rate
  2. Calculate the​ town's labor force participation rate
  3. Calculate Suppose that 10,000 people decide to join the labor force. Of these 10,000 people 7,800 find jobs. What is the unemployment rate at the end of 2018?
Answer #1

At beginning of 2018

No. of employed people= 152,000

No of unemployed people= 8000

Population not in labour force= 40,000

a) Town's unemployment rate= No.of people unemployed/ Total working population

= 8000/160000*100

= 0.05*100

= 5%

b) Towns labour force participation rate

=Employed people + unemployed people / working age non-institunalized population

= 152,000+8000/152000+8000+40000*100

= 160000/200000*100

= 0.8*100

= 80%

At the end of 2018, suppose 10000 people joined the labour force

So total labour force increases to 152000+8000+10000=170000

Of these 7800 find jobs so unemployed here is 2200, so total unemployed increases to 8000+2200=10200

So unemployment rate at the end of 2018

=10200/170000*100

=0.06*100

=6%